Please use the big M method to solve the following linear program. Write down all tableau, note basic variables and nonbasic variables. Use slack and artificial variables. Construct your tableau iterations using the standard form of the program. For example first line z+2x1-2x2+2x3=0. If possible, STATE THE OPTIMAL SOLUTION AND THE OPTIMAL VALUE. Otherwise state why you cannot find them.

Consider the following linear program: 2x3 max z= –2x1 + s.t. + -x1 21 > 0, 2x2 - 2x2 + + x2 > 0, x3 > 03 > -1 0
Answer #1

solation or max2 = -289 +2212 - 20g st 24 +242 + x3 = 4 +29+ 8g 2nd see, 2d, 292,0 Standard for max2 = -229,7242-200 g tosi-mPROBLEM: -M -M C -2 2 -2 0 Var | X1 X2 X3 S2 A1 2 2 1 0 A2 -1 0 1 -1 Z -M -2M-2M M C-1-2+M 2+2M 2+2 -M M A1 1 0 M 0 M A2 0 4